OpenStudy (anonymous):

Ill Give Medals Nellie is analyzing a quadratic function f(x) and a linear function g(x). Will they intersect? x g(x) 1 5 2 10 3 15

OpenStudy (anonymous):

Yes, at positive x-coordinates Yes, at negative x-coordinates Yes, at negative and positive x-coordinates No, they will not intersect

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i need i dont know this stuff and i cant find it in my lesson

OpenStudy (anonymous):

I need Help

OpenStudy (anonymous):

I think you should try plotting the graph of g(x) and if there's a point that will pass through the graph of f(x), then they have an intersection.

OpenStudy (anonymous):

i did and i think its B

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

gimme a sec. i'll take a look

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Its either C or B

OpenStudy (anonymous):

hello @Data_LG2

OpenStudy (anonymous):

Can you help @CynfulBaddie

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

can you help @Hotchellerae21

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

hello i think its C now looking at the problem @Hotchellerae21 @Data_LG2

OpenStudy (anonymous):

i think it is C too i don't know

OpenStudy (anonymous):

it could also Be B

OpenStudy (anonymous):

can you help me @amistre64

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

@data

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

ill give you all medals

OpenStudy (anonymous):

If you help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes what?

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Can you help @dan815

OpenStudy (dan815):

Figure out what the equation of that quadratic function is

OpenStudy (anonymous):

thanks @dan815

OpenStudy (anonymous):

i pretty sure its B

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

y=ax^2+bx+c are x=0, y=2 so 2=a(0^2)+b(0)+c therefore c=2 y=ax^2+bx+2 Now, sub in another couple points 4=a(2^2)+b(2)+2 4=a(-2)^2+b(-2)+2 ---------------- 4=4a+2b+2 4=4a-2b+2 ------------- ^---- solving these 2 equations gives us a=1/2 and b=0 y=1/2x^2+2 <-- is our equation

OpenStudy (anonymous):

so is it B Because im pretty sure it is

OpenStudy (anonymous):

@dan815 is it B

OpenStudy (dan815):

y=mx+b is the equation of a line, i want to rush a bit so y=5x is the eqn of your line lets see if theres an intersection y=5x y=1/2x^2+2 5x=1/2x^2+2 10x=x^2+4 x^2-10x+4=0 \[+or-b+\frac{\sqrt{(b^2-4ac)}}{2a}\] under the radical b^2-4ac= 100-(4*1*4) <--- isnon negative so, yes there will be intersection points, infact 2 intersection points

OpenStudy (dan815):